OpenStudy (anonymous):

If f(x) = x + 5, find f(-9).

OpenStudy (anonymous):

\[f(x) = x + 5\] \[f(-9) = (-9) + 5\] \[f(-9) = -4\]

OpenStudy (precal):

f(-9)=-4 is function notation it also, represents (-9,-4) this is just coordinate notation